A investor invest a total of $2000 in two mutual funds. One earned a profit of 4% while the other earned 2%. His total profit was 56$ How much money did he put in each one? I am having trouble finding a formula for this problem
Answer by stanbon(75887)

A investor invest a total of $2000 in two mutual funds. One earned a profit of 4% while the other earned 2%. His total profit was 56$ How much money did he put in each one?
---------------
Quantity Equation: x + y = 2000
Interest Equation: 4x +2y = 5600
------------------------------------
Modify:
2x + 2y = 4000
4x + 2y = 5600
----
Subtract and solve for "x":
2x = 1600
----
x = $800 (amt. invested at 4%)
y = $1200 (amt. invested at 2%)
=================
Cheers,
Stan H.